There are two numbers that add up to 132. One number is the square root of another, what are those two numbers?

Answer by Theo(13342)

x + sqrt(x) = 132
subtract x from both sides of the equation to get:
sqrt(x) = 132 - x
square both sides of the equation to get:
x = 17424 - 264x + x^2
subtract x from both sides of the equation to get:
0 = 17424 - 265x + x^2
use the quadratic formula to solve for x.
you will get:
x1 = 144
x2 = 121
plug these values of x into the original equation and you will find that the solution to the problem has to be x = 121.
121 + sqrt(121) = 132 becomes:
121 + 11 = 132 which becomes:
132 = 132
solution is confirmed.

moving the x from the left side of the equation to the right side of the equation allowed you to square both sides of the equation which eliminated the sqrt(x) and allowed you to solve it easily.